dfrancis:

With a minimum salary of 1350 per month plus commission of 4.45 % of her total sales. What are Lorenas total sales for a month in which her gross pay is 2501.22

SmokeyBrown:

Hi and welcome to QuestionCove! We can model Lorena's pay as an equation based on the information we're given: \[Gross pay = 1350 + 0.0445*Sales\] For this month, we know her Gross pay is equal to 2501.22, so we can put that into the equation: \[2501.22 = 1350 + 0.0445*Sales\] Next, let us try to solve for Sales using algebra. First, we'll subtract 1350 from both sides of the equation. The result of this operation would be: \[1151.22 = 0.0445*Sales\] Then, we can divide both sides of the equation in order to get Sales by itself on the right side. This will allow us to solve for "Sales" and calculate Lorena's sales for that month. We get: \[1151.22/0.0445 = Sales\] So, what would be the total value of Sales at the end of these operations?